OpenStudy (anonymous):

Find the maclaurin series for f(x) when f(x) = ln(1-x^4)

OpenStudy (anonymous):

a good gimmick would be to take the derivative, get that power series, then integrate term by term

OpenStudy (anonymous):

that way you don't have to keep taking successive derivatives

OpenStudy (anonymous):

I dont know how to do that

OpenStudy (anonymous):

just take derivative of ln(1-x^4)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh wait, maybe that is not going to work here

OpenStudy (anonymous):

damn ok then you have to do it by hand

OpenStudy (anonymous):

\(f(0)=0\) so the constant is 0

OpenStudy (anonymous):

\[f'(x)=\frac{4x^3}{1-x^4}\] so \[f'(0)=0\] as well. hmmm

OpenStudy (anonymous):

actually it is \[f'(x)=\frac{-4x^3}{1-x^4}\]

OpenStudy (anonymous):

ok I got derivative but im not sure what to do next, the answer has summation in front of it like \[f(x)=\sum_{n=1}^{infinity}\]

OpenStudy (anonymous):

and then asks for your answer

OpenStudy (anonymous):

well this is going to take a lot of work i think

OpenStudy (anonymous):

you are going to get zeros until the \(x^4\) term

OpenStudy (anonymous):

I dont mind I just dont know what the steps I need to take to get there

OpenStudy (anonymous):

okay i was trying to think of a better way to do it, maybe for example writing it as \[\log(1-x^4)=\log(1+x^2)+\log(1-x)+\log(1+x)\]

OpenStudy (anonymous):

you have to keep taking successive derivatives, and evaluate them at 0 those are your coefficients

OpenStudy (anonymous):

well not exactly, it is \(\frac{f^{n}(0)}{n!}x^n\)

OpenStudy (anonymous):

i can't come up with a better way of doing it than by hand

OpenStudy (anonymous):

lol its ok I gave up, ill just ask my professor tomorrow, I just didnt go to class today

OpenStudy (anonymous):

thx anyway

OpenStudy (anonymous):

which is going to suck because the derivatives get nastier and nastier first 4 will give 0, so the first non zero term is the \(x^4\) term

OpenStudy (anonymous):

you might try using \[\log(1-x^4)=\log(1+x^2)+\log(1-x)+\log(1+x)\] and expanding each term the mclauren series for the last two are well known

OpenStudy (anonymous):

\[\log(1-x)=-\sum_{n=1}^{\infty}\frac{x^n}{n}\]

OpenStudy (anonymous):

\[\log(1+x)=\sum_{n=1}^{\infty}(-1)^{n+1}\frac{x^n}{n}\]